OpenStudy (anonymous):

You wad up a piece of paper and throw it into the wastebasket. How far will your piece of paper travel if you throw it with a velocity of 1.4 m/s at an angle of 36? (Assume no air resistance and that the paper lands at the same height you threw it from.) A. 0.77 m B. 0.48 m C. 0.19 m D. 0.25 m

OpenStudy (anonymous):

There's a useful formula for projectiles (with no air resistance, etc) that is pretty simple to demonstrate: Horizontal distance of the projectile = \[Distance = \frac{ V_{0}^{2} }{ g}2\sin \theta \cos \theta\] That'll help :D